Let Px be a quadratic polynomial such that for distinct realms alpha and beta p-alpha =alpha p-beta=beta show that alpha and beta are roots of p(p(x)) -x=0
1
Expert's answer
2020-04-20T18:35:46-0400

Let "P(x)" be a quadratic polynomial such that for distinct realms "\alpha" and "\beta" "P(-\alpha) =\alpha, P(-\beta)=\beta" show that alpha and beta are roots of "P(P(x)) -x=0"


Let "P(x)=a_0x^2+a_1x+a_2, a_0,a_1,a_2\in R, a_0\neq0"

"P(-\alpha)=a_0\alpha^2-a_1\alpha+a_2=\alpha\\
a_0\alpha^2+a_2=\alpha+a_1\alpha"

"P(P(x))-x=a_0(a_0x^2+a_1x+a_2)^2+\\
+a_1(a_0x^2+a_1x+a_2)+a_2-x"

Then

"P(P(\alpha))-\alpha=a_0(a_0\alpha^2+a_1\alpha+a_2)^2+\\
+a_1(a_0\alpha^2+a_1\alpha+a_2)+a_2-\alpha=\\
=a_0(\alpha+2a_1\alpha)^2+a_1(\alpha+2a_1\alpha)+a_2-\alpha=\\
=a_0\alpha^2+4a_0a_1\alpha^2+4a_0a_1^2\alpha^2+a_1\alpha+2a_1^2\alpha+\\
+a_2-\alpha=\alpha+a_1\alpha+4a_0a_1\alpha^2+4a_0a_1^2\alpha^2+\\
+a_1\alpha+2a_1^2\alpha-\alpha=\\
=4a_0a_1\alpha^2+4a_0a_1^2\alpha^2+\\
+2a_1\alpha+2a_1^2\alpha=4a_0a_1\alpha^2(1+a_1)+\\
+2a_1\alpha(1+a_1)=\\=2a_1\alpha(1+a_1)(2a_0\alpha+1)=0"

becase "a_1=-1"

"P(1)=a_0+a_1+a_2" - the sum of the coefficients of the polynomial

"a_0-a_1+a_2-1=0"


Similarly "P(P(\beta))-\beta=0"

or

"P(P(\beta))-\beta=P(-(-P(\beta)))-\beta=-P(\beta)-\beta=\\
-P(-(-\beta))-\beta=-(-\beta)-\beta=\beta-\beta=0"